If 0.590% solution of Na²CO3 will be used to precipitate Ca+² from solution, determine the amount of solution, in grams, needed to precipitate 5.00×10²mL of 0.01120 M Ca+²?

1
Expert's answer
2022-03-13T15:16:38-0400

CaCO3(s) = Ca2+(aq) + CO32-(aq)

Ksp = 4.8 × 10-9

n(Ca2+) = 0.01120 mol/l × 5.00×102×10-3 l = 0.0056 mol

0.590 g Na2CO3 in 100 g solution

n(Na2CO3) = m(Na2CO3) / Mr(Na2CO3) = 0.590 g / 106 g/mol = 0.005566 mol

V(solution) = 100 g / 1 g/ml = 100 ml = 0.1 l

M = 0.005566 mol / 0.1 l = 0.05566 mol/l

Ksp = [Ca2+] × [CO32-]

4.8×10-9 = 0.0056 mol × [CO32-]

[CO32-] = 4.8×10-9 / 0.0056 = 8.57×10-7 mol

V(solution) = 8.57×10-7 mol / 0.05566 mol/l = 1.54×10-5 l = 0.0154 ml

m(solution) = 0.0154 ml × 1 g/ml = 0.0154 g